OpenStudy (anonymous):

Meshia spent $14.50 on a total of 10 items. She bought soft drinks for $1.25 each and popcorn for $1.75 each. How many soft drinks did she buy?

OpenStudy (amistre64):

i got these worked down to one formula: amount @ y = t(z-x)/(y-x) amount @ 1.25 = (14.50-10(1.75))/(1.25-1.75) perhaps?

OpenStudy (amistre64):

if i did it right; I get 6 at 1.25; gotta double check tho

OpenStudy (anonymous):

ok.thanks

OpenStudy (amistre64):

6(1.25) + 4(1.75) = 14.50 7.50 + 7.00 = 14.50 .. looks good to me :)

OpenStudy (anonymous):

right

OpenStudy (amistre64):

actually 2 formulas amount @ y = t(z-x)/(y-x); which this doesnt call for and amount @ y = (z-tx)/(y-x) ; which this one uses

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

to work out the details tho; you wanna create 2 equations a + b = t (for total) ax + by = z (where z is the amount given), this one we need now we can use elimination on these a + b = t <-- multiply by -x ax + by = z -ax -bx = -tx ax + by = z -------------- ; and add them up b(y-x) = z - tx ; and solve for b b = (z-tx)/(y-x) ...